OpenStudy (anonymous):

find three consecutive odd integers whose sum is -33?

Parth (parthkohli):

x + x + 1 + x + 2 = -33 3x + 3 = -33 3x = -33-3 = -36 x = -36/3 x = -12 So, answer is: -12,-11,-10

OpenStudy (anonymous):

Easiert way would be this: Take the integers as x-1, x and x+1 Their sum is 33. You should be able to work out the algebra.